# Large Deviations Principle for Isoperimetry and Its Equivalence to Nonlinear Log-Sobolev Inequalities

### v2

@@ 1.1.2. Companion of Various Constants
We apply the equivalence in Theorem 1 to compare the log-Sobolev constant, the isoperimetric constant, and constants for various other inequalities. To this end, we define the following various constants.
@@
  • $K_{T}^{+}$ denotes the optimal constant $K$ in the transport inequality that there is some constant $D_{0}$ such that for all
@@
  • $K_{C}^{+}$ denotes the optimal constant $K$ in the asymptotic dimension-free Gaussian concentration inequality that there is some constant $r_{0}$ such that for all $n\geq 1$ and $r\geq r_{0}$, $\nu_{n}((A^{r})^{c})\leq e^{-\frac{Kr^{2}}{2}}$ for all $A$ such that $\nu_{n}(A)\geq 1/2$.
@@
  • $K_{IS}^{-}$ denotes the constant in the dimension-free (also infinite dimensional) isoperimetric inequality
@@
We now compare all optimal constants defined above.
@@
Assume that (1.5) holds. The following hold.
@@
    (1.13) $K_{CD}\overset{\mathrm{(a)}}{\leq}\left(\frac{1+K_{CD}/K_{C}^{+}}{2}\right)^{2}K_{C}^{+}\overset{\mathrm{(b)}}{\leq}$ $K_{IS}^{-}\overset{\mathrm{(c)}}{\leq}K_{IS}\overset{\mathrm{(d)}}{=}K_{LS}\overset{\mathrm{(e)}}{=}K_{HC}$
@@
    $K_{LS}\overset{\mathrm{(j)}}{\leq}K_{LS}^{+}\overset{\mathrm{(k)}}{\leq}K_{T}^{+}\wedge K_{IS}^{+},$
    where $x\wedge y=\min\{x,y\}$, and the equality (d) holds under the additional assumption (1.8).
  2. (2)
    Assume further (1.8). It holds that $K_{C}^{+}\leq 4K_{IS}^{-}$, i.e., all the quantities in (1.13) except for $K_{CD}$ differ at most 44 times. In particular, $\Theta$ is sandwiched between two rays $\alpha\mapsto K_{LS}\alpha$ and $\alpha\mapsto K_{LS}^{+}\alpha$ with
    $K_{LS}^{+}/4\leq K_{LS}\leq K_{LS}^{+}.$
@@
For Statement 1, Inequality (a) is trivial, by noting that $0\leq K_{CD}/K_{C}^{+}\leq 1$. Inequality (c) is due to the fact that $n\mapsto I_{n}(a)$ is nonincreasing. Equality (d) is implied by Theorem 1. Equality (e) is a classic result of Gross [24]. The inequality $K_{LS}\leq K_{T}$ (or equivalently, Inequality (f)) was proven by Otto and Villani [38]. Equality (g) is also well known; see e.g. [47, Remark 22.23]. Inequalities (h) and (j) follow by definition. Equality (i) follows similarly to (g). The part $K_{LS}^{+}\leq K_{T}^{+}$ in Inequality (k) follows similarly to $K_{LS}\leq K_{T}$. The part $K_{LS}^{+}\leq K_{IS}^{+}$ in Inequality (k) follows by Theorem 1, with a detailed proof given in Section 6. So, it only remains to prove Inequality (b). This is done in Section 6. In addition, the inequality $\left(\frac{1+K_{CD}/K_{C}}{2}\right)^{2}K_{C}\leq K_{LS}$ was already known; see Theorem 22.21 of [47]. However, we do not know if this implies the inequality (b) since we do not know if $K_{IS}^{-}=K_{IS}$ holds. Statement 2 is a consequence of Statement 1. Statement 3 is proven in Section 6.
As a consequence of Theorem 2, the following corollary is obtained.
Corollary 1.
If $K_{CD}$ is equal to any one of other quantities in (1.13), then all inequalities (a), (b), (c), (f), and (h) in (1.13) become equalities.
Without the assumption (1.8), we have the following weaker result, as a consequence of Theorem 2.
@@
Assume that (1.5) holds. Then, it holds that for any $\alpha>0$,
(1.14) $\sqrt{\frac{K_{C}^{+}}{4}\alpha}\leq\underline{\Lambda}(\alpha)\leq\overline{\Lambda}(\alpha)\leq\sqrt{\breve{\Theta}(\alpha)}$}.
Let $I_{\infty}$ be the isoperimetric profile defined for the infinite dimensional product space $\mathcal{M}^{\infty}$ equipped with the product topology and the product measure $\nu_{\infty}:=\nu^{\otimes\infty}$, but the enlargement of a set is defined under $d_{\infty}(x,y)=\sqrt{\sum_{i=1}^{\infty}d(x_{i},y_{i})^{2}}$ (although it might diverse to infinity for two arbitrary points). Then, it is obvious that $I_{\inf}(a)\geq I_{\infty}(a)$. In the following, we show that this is actually an equality. That is, the isoperimetric profile $I_{\infty}$ on the infinite dimensional space is the pointwise limit of the isoperimetric profiles on finite dimensional spaces.
@@
Assume that (1.5) holds. Then, $I_{\inf}(a)=I_{\infty}(a)$ for all $a\in[0,1]$.
The proof is provided in Section 7.
@@ 1.2. Open Problems
Assume that (1.5) holds. Is it true that $K_{IS}^{-}=K_{LS}$?
A more challenging problem is to determine the function $I_{\inf}$ (or equivalently, $I_{\infty}$). This problem is a commonly recognized open one.
@@
Assume that (1.5) holds. What is the explicit expression of $I_{\inf}$?
Another interesting topic is to remove the assumption (1.8) from Theorem 1. We believe Theorem 1 still holds even without this assumption. To this end, we propose the following conjecture.
@@
Let $\nu_{1}=e^{-V_{1}}\mathrm{vol}_{1},\nu_{2}=e^{-V_{2}}\mathrm{vol}_{2}$ with $V_{1}\in C^{2}(\mathcal{M}_{1}),V_{2}\in C^{2}(\mathcal{M}_{2})$ be probability measures respectively on Riemannian manifolds $\mathcal{M}_{1},\mathcal{M}_{2}$. Assume both $V_{1}$ and $V_{2}$ satisfy (1.5). Then,
@@
This conjecture can be equivalently formulated as the plane version, as explained in the following. Any probability measure $\nu$ satisfying (1.5) admits a concave isoperimetric profile $I_{\nu}$. For every such probability measure $\nu$, one can find a unique symmetric (w.r.t. $x=0$) log-concave probability measure $\tilde{\nu}$ on $\mathbb{R}$ such that $I_{\tilde{\nu}}=I_{\nu}$ [10]. Explicitly, the cdf $F$ of $\tilde{\nu}$ is exactly the inverse of the function
@@
which will induces $I_{\tilde{\nu}}=f\circ F^{-1}=I_{\nu}$, with $f=F^{\prime}$ denoting the pdf of $\nu$. Let $\nu_{1}=e^{-V_{1}}\mathrm{vol},\nu_{2}=e^{-V_{2}}\mathrm{vol}$ with $V_{1}\in C^{2}(\mathcal{M}_{1}),V_{2}\in C^{2}(\mathcal{M}_{2})$ be probability measures respectively on Riemannian manifolds $\mathcal{M}_{1},\mathcal{M}_{2}$. Assume both $V_{1}$ and $V_{2}$ satisfy (1.5). Let $\tilde{\nu}_{1},\tilde{\nu}_{2}$ be the symmetric log-concave probability measures on $\mathbb{R}$ induced by $\nu_{1},\nu_{2}$ respectively such that $I_{\tilde{\nu}_{i}}=I_{\nu_{i}},i=1,2$. Denote $F_{i},f_{i}$ as the cdf and pdf of $\tilde{\nu}_{i}$, $i=1,2$. Define
@@
where the first infimum is taken over all smooth functions $\sigma:\mathcal{M}_{1}\to[0,1]$ such that $\int_{\mathcal{M}_{1}}\sigma(x){\text{\rm d}}\nu_{1}(x)=a$, the second infimum is taken over all smooth functions $\varphi:[0,1]\to[0,1]$ such that $\int_{0}^{1}\varphi(t){\text{\rm d}}t=a$, and the third infimum is taken over all smooth functions $g:\mathbb{R}\to\mathbb{R}$ such that $\int_{\mathbb{R}}F_{2}(g(x))f_{1}(x){\text{\rm d}}x=a$. The following proposition shows that the isoperimetric profile of $\nu_{1}\otimes\nu_{2}$ will decrease if we replace one or both of $\nu_{1},\nu_{2}$ to the corresponding $\tilde{\nu}_{1},\tilde{\nu}_{2}$. This proposition is an extension of the Gaussian version due to Bobkov [9] to Riemannian manifolds. It might be already known, but we have not found it in the literature. So, we provide a proof in Section 8 for completeness.
@@
holds for all smooth functions $\sigma_{i}:\mathcal{M}_{i}\to[0,1]$. Then, by an induction method and applying Proposition 2, $I_{\nu_{1}\otimes\nu_{2}\otimes...\otimes\nu_{i}}\geq\Upsilon_{i},i\in\{1,2,...,n\}$.
@@ 1.3. Related Works
In the one-dimensional setting, Bobkov [7] established that half-lines are extremal sets in the isoperimetric inequality for log-concave probability measures. Extending to higher-dimensional structures, Bobkov and Houdré [11] further explored the isoperimetric constant and log-Sobolev inequalities in product spaces, laying a foundation for multi-dimensional isoperimetric analysis. The optimality of half-spaces— a core topic in isoperimetry— has been systematically studied by multiple scholars: Sudakov and Tsirel’son [43] initiated early investigations, Bobkov [7] deepened this topic by addressing the isoperimetric problem for log-concave product measures on $\mathbb{R}^{n}$ equipped with the uniform distance, and explicitly provided necessary and sufficient conditions for standard half-spaces to be extremal, and Barthe [5] investigated the extremal properties of central half-spaces (with measure $1/2$) for product probability measures on Riemannian manifolds equipped with the geodesic distance. It should be noting that the isoperimetric problem under the geodesic distance and the one under the uniform distance are significantly different. When equipped with the uniform distance, the asymptotics of the isoperimetric profile has been characterized by Bobkov in [8]. However, for the geodesic distance as in our setup, the asymptotics is still widely unknown except for special cases, e.g., Gaussian measures. For a comprehensive overview of isoperimetry in product spaces, Roberto [41] offers a critical survey that synthesizes key results in the field. It should be noted that characterizing the isoperimetric profile is widely open.
@@
Concentration of measure has natural connections to isoperimetric inequalities. In fact, it can be seen as a variant version of isoperimetric problem concerning on sets with probability measure $1/2$ in which the boundary of a set is set to the $r$-neighbourhood of the set with $r>0$. A landmark methodological contribution in concentration of measure was made by Marton [31], who was the first to introduce information-theoretic techniques (especially transport-entropy inequalities) to the study of concentration of measure, yielding an elegant and concise proof for a dimension-free bound on concentration of measure. Talagrand extended this idea by developing a new transport-entropy inequality, adapting it to Gaussian measures and the Euclidean metric [45]; this argument has since become a standard reference, featured in prominent textbooks [27, 40, 46]. Gozlan and Léonard further refined this approach by replacing the “linear” transport-entropy inequality in Marton’s argument with a “nonlinear” version, obtaining a dimension-free sharp bound on concentration of measure—exponentially tight, as the exponent of the bound is asymptotically attained [22]. Prior to Gozlan and Léonard’s work, Alon et al. [3] are the first to characterize the sharp convergence exponent for concentration of measure, but limited to finite spaces. Gozlan [21] also used Marton’s framework to prove the equivalence between the Gaussian bound and Talagrand’s transport-entropy inequality. Additionally, Dembo [17] proposed a new class of transport-entropy inequalities, which he used to recover and generalize several of Talagrand’s key results [44].
Ahlswede and Zhang [2] and Yu [48] contributed to the study of isoperimetry with thick boundaries for exponential small sets. Ahlswede and Zhang [2] focused on finite spaces for which a key tool called the inherently typical subset lemma was developed by them. Yu [48] generalized their result to Polish spaces (complete, separable metric spaces), expanding the problem’s applicability, especially to the setting considered in this paper (i.e., the isoperimetric problem on a manifold with density). This plays a core role in our proof in the present paper. Furthermore, E. Milman [32] investigated the equivalence between isoperimetric problems with thick and thin boundaries under specific conditions. There are many existing works investigating isoperimetric inequalities or concentration of measure on various graphs, e.g., [30, 1, 12, 19].
@@
Ledoux [26, 28] proposed a semigroup approach to derive isoperimetric inequalities from hypercontractivity (which is equivalent to log-Sobolev inequalities). However, the constant obtained via this method is suboptimal. In contrast, the constant in our work is asymptotically optimal. Otto and Villani [38] established that log-Sobolev inequalities imply transport-entropy inequalities, bridging log-Sobolev theory with concentration of measure. Building on this, Ledoux [29] (alongside Milman [32]) used a semigroup approach to derive isoperimetric inequalities from concentration inequalities (equivalently, transport-entropy inequalities). Similar to his earlier work, Ledoux’s constant here is suboptimal, whereas our constant achieves asymptotic optimality. The concept of nonlinear Sobolev inequalities is due to Polyanskiy and Samorodnitsky [39], and they showed a pseudo tensorization property for it. Their work opened a new direction in Sobolev theory, allowing for more flexible and problem-specific inequality formulations. Although nonlinear Sobolev inequalities were originally proposed in [39] not for purpose of studying isoperimetry, in the present paper we observe that there is indeed an exact equivalence between them.
1.4. Proof Strategy and Paper Organization

### v3

@@ 1.1.2. Companion of Various Constants
While Theorem 1 is primarily tailored to subsets with probability measures decaying exponentially in $n$, it admits application to bounding the isoperimetric profile for subsets with probability measures decaying arbitrarily slowly. In other words, Theorem 1 can be used to quantitatively compare the log-Sobolev constant, the isoperimetric constant, and other constants for various inequalities. To this end, we define the following constants.
@@
  • $K_{T}^{+}$ denotes the optimal constant $K$ in the transport inequality that there is some constant $D_{0}\geq 0$ such that for all
@@
  • $K_{C}^{+}$ denotes the optimal constant $K$ in the asymptotic Gaussian concentration inequality that there is some constant $r_{0}\geq 0$ such that for all $n\geq 1$ and $r\geq\sqrt{n}r_{0}$, $\nu_{n}((A^{r})^{c})\leq e^{-\frac{Kr^{2}}{2}}$ for all $A$ such that $\nu_{n}(A)\geq 1/2$.
@@
  • $K_{IS}^{-}$ denotes the constant in the dimension-free (also infinite-dimensional) isoperimetric inequality
@@
From the definitions, the constants $K_{IS}$, $K_{IS}^{+}$, and $K_{IS}^{-}$ are the asymptotic isoperimetric ratios $\frac{I_{n}(a)}{a\sqrt{2\ln\frac{1}{a}}}$ respectively for exponentially, superexponentially, and subexponentially small sequences of volumes. We now compare all optimal constants defined above.
@@
Under Assumption 1 (i.e., $K_{CD}\geq 0$), the following hold.
@@
    (1.14) $K_{CD}\overset{\mathrm{(a)}}{\leq}\left(\frac{1+K_{CD}/K_{C}}{2}\right)^{2}K_{C}\overset{\mathrm{(b)}}{\leq}$ $K_{IS}^{-}\overset{\mathrm{(c)}}{\leq}K_{IS}\overset{\mathrm{(d)}}{\leq}K_{LS}\overset{\mathrm{(e)}}{=}K_{HC}$
@@
    $K_{LS}\overset{\mathrm{(j)}}{\leq}K_{LS}^{+}\overset{\mathrm{(k)}}{\leq}K_{T}^{+},$
    where $x\wedge y=\min\{x,y\}$. Under the additional Assumption 2, the equality in (d) holds and moreover,
    $K_{LS}^{+}\overset{\mathrm{(l)}}{\leq}K_{IS}^{+}.$
    In addition, if $K_{CD}$ is equal to any one of $K_{IS}^{-},K_{IS},K_{LS},K_{HC},K_{T},K_{C}$, then all the inequalities in (a)–(g) of (1.14) become equalities.
@@
For Statement 1, Inequality (a) is trivial, by noting that $0\leq K_{CD}/K_{C}\leq 1$. Inequality (c) is due to the fact that $n\mapsto I_{n}(a)$ is nonincreasing. Equality (d) is implied by Theorem 1. Equality (e) is a classic result of Gross [33]. The inequality $K_{LS}\leq K_{T}$ (or equivalently, Inequality (f)) was proven by Otto and Villani [54]. Equality (g) is also well known; see e.g. [65, Remark 22.23]. Inequalities (h) and (j) follow by definition. Equality (i) follows similarly to (g). Inequality (k) follows similarly to $K_{LS}\leq K_{T}$. Inequality (l) follows by Theorem 1, with a detailed proof given in Appendix A. So, it only remains to prove Inequality (b). This is done in Appendix A. In addition, a weaker version of (b), $\left(\frac{1+K_{CD}/K_{C}}{2}\right)^{2}K_{C}\leq K_{LS}$, was already known; see Theorem 22.21 of [65]. However, we do not know if this is equivalent to the inequality (b) since we do not know if $K_{IS}^{-}=K_{IS}$ holds. Statement 2 is proven in Appendix A.
Example 3 (Hypercube).
We continue considering the hypercube described in Example 2. Statements in Example 2 imply that for this case,
$K_{CD}=0,\;K_{IS}=K_{LS}=K_{HC}=K_{T}=K_{C}=\pi^{2},\;K_{LS}^{+}=K_{C}^{+}=K_{T}^{+}=K_{IS}^{+}=\infty,$
where the density $1+\epsilon\cos\pi x$ asymptotically attains the constants $K_{LS}$ and $K_{T}$ as $\epsilon\to 0$. Moreover, it holds that $\sqrt{2\pi}\leq K_{IS}^{-}\leq\pi^{2}$; the lower bound will be explaind in Example 4. However, the exact value of $K_{IS}^{-}$ is still unknown.
Without Assumption 2, we have the following weaker result, as a consequence of Theorem 2.
@@
Under Assumption 1, it holds that for any $\alpha>0$,
(1.15) $\sqrt{\frac{K_{LS}}{4}\alpha}\leq\sqrt{\frac{K_{T}}{4}\alpha}\leq\underline{\Lambda}(\alpha)\leq\overline{\Lambda}(\alpha)\leq\sqrt{\breve{\Theta}(\alpha)}$}.
1.1.3. Central Limit Regime
Let $I_{\infty}$ be the isoperimetric profile defined for the infinite dimensional product space $\mathcal{M}^{\infty}$ equipped with the product topology and the product measure $\nu_{\infty}:=\nu^{\otimes\infty}$, but the enlargement of a set is defined under $d_{\infty}(x,y)=\sqrt{\sum_{i=1}^{\infty}d(x_{i},y_{i})^{2}}$ (although it might diverse to infinity for two arbitrary points). Then, it is obvious that $I_{\inf}(a)\geq I_{\infty}(a)$. In the following, we show that this is actually an equality. That is, the isoperimetric profile $I_{\infty}$ on the infinite dimensional space is the pointwise limit of the isoperimetric profiles on finite dimensional spaces. The proof is provided in Appendix B.
@@ 1.1.3. Central Limit Regime
Assume that $I_{\inf}$ is continuous (e.g., Assumption 1 holds). Then, $I_{\inf}(a)=I_{\infty}(a)$ for all $a\in[0,1]$.
A well-known lower bound on $I_{\inf}$ is that
(1.16) $I_{\inf}(a)\geq K_{0}\cdot I_{\mathrm{G}}(a),\;\forall a\in[0,1],$
where $I_{\mathrm{G}}$ is the Gaussian isoperimetric profile and $K_{0}:=\inf_{a\in(0,1)}\frac{I_{\nu}(a)}{I_{\mathrm{G}}(a)}$ [13, 8]. An upper bound can be obtained by considering the $\mathbf{f}$-coordinate half-space given in (1.8), as shown in the following theorem. The proof is based on the coarea formula and the central limit theorem, which is given in Appendix C.
Theorem 5 (Upper Bound for Isoperimetry in Central Limit Regime).
Assume that $I_{\inf}$ is continuous (e.g., Assumption 1 holds). Then, for any $a\in[0,1]$,
(1.17) $I_{\inf}(a)\leq\sqrt{K_{P}}\cdot I_{\mathrm{G}}(a),$
where $K_{P}$ is the spectral gap (the reciprocal of the Poincaré constant) for $\nu$, i.e., the optimal constant in
(1.18) $K_{P}\cdot\mathrm{Var}_{\nu}(f)\leq\mathbb{E}_{\nu}\left[\|\nabla f\|^{2}\right]$
for all $f$ in the Sobolev space $W^{1,2}(\mathcal{M})$.
As a related result, the Poincaré constant and the Cheeger constant are known to be closely related; see e.g., [35]. The Cheeger constant characterizes the isoperimetric behavior exclusively for large subsets, whereas the inequality in (1.17) captures isoperimetry for all subsets without restriction.
The upper bound $\sqrt{K_{P}}I_{\mathrm{G}}(a)$ in (1.17) can be used to bound $K_{IS}^{-}$, namely, obtaining that $K_{IS}^{-}\leq K_{P}$. However, this bound is worse than the upper bound $K_{LS}$, since in general $K_{LS}\leq K_{P}$. Note that by perturbation analysis, one can see that
$K_{P}=\liminf_{\alpha\to 0}\frac{\Theta(\alpha)}{2\alpha}.$
Thus, $K_{LS}=K_{P}$ if $\Theta(\alpha)\geq 2K_{P}\alpha$ (e.g., if $\Theta$ is convex), and otherwise, $K_{LS}<K_{P}$ (e.g., for exponential measures [65, Example 21.19]). For the latter case, the upper bound $\sqrt{K_{P}}I_{\mathrm{G}}(a)$ is not tight for sufficiently small $a$. Combining the best known lower bound in (1.16), the upper bound in (1.17), and the trivial upper bound $I_{1}$ yields
$K_{0}I_{\mathrm{G}}(a)\leq I_{\inf}(a)\leq I_{1}(a)\wedge\sqrt{K_{P}}I_{\mathrm{G}}(a).$
Example 4 (Hypercube).
We continue considering the hypercube described in Example 2. Applying the inequalities above to the hypercube $(0,1)^{n}$ yields $\sqrt{2\pi}I_{\mathrm{G}}(a)\leq I_{\inf}(a)\leq 1\wedge\pi I_{\mathrm{G}}(a)$. As shown in [29], the lower bound is not tight.
@@ 1.2. Open Problems
Under Assumption 1, is it true that $K_{IS}^{-}=K_{LS}$?
A more challenging problem is to exactly determine the function $I_{\inf}$ (or equivalently, $I_{\infty}$). This problem is a commonly recognized open one.
@@
Under Assumption 1, what is the explicit expression of $I_{\inf}$?
Another interesting topic is to remove Assumption 2 from Theorem 1. To this end, we propose the following conjecture.
@@
Let $\nu_{1}=e^{-V_{1}}\mathrm{vol}_{1},\nu_{2}=e^{-V_{2}}\mathrm{vol}_{2}$ with $V_{1}\in C^{2}(\mathcal{M}_{1}),V_{2}\in C^{2}(\mathcal{M}_{2})$ be probability measures respectively on Riemannian manifolds $\mathcal{M}_{1},\mathcal{M}_{2}$. Assume both $V_{1}$ and $V_{2}$ satisfy Assumption 1. Then,
@@
This conjecture can be equivalently formulated as the plane version, as explained in the following. Any probability measure $\nu$ satisfying Assumption 1 admits a concave isoperimetric profile $I_{\nu}$. For every such probability measure $\nu$, one can find a unique symmetric (w.r.t. $x=0$) log-concave probability measure $\tilde{\nu}$ on $\mathbb{R}$ such that $I_{\tilde{\nu}}=I_{\nu}$ [14]. Explicitly, the cdf $F$ of $\tilde{\nu}$ is exactly the inverse of the function
@@
which will induces $I_{\tilde{\nu}}=f\circ F^{-1}=I_{\nu}$, with $f=F^{\prime}$ denoting the pdf of $\nu$. Let $\nu_{1}=e^{-V_{1}}\mathrm{vol},\nu_{2}=e^{-V_{2}}\mathrm{vol}$ with $V_{1}\in C^{2}(\mathcal{M}_{1}),V_{2}\in C^{2}(\mathcal{M}_{2})$ be probability measures respectively on Riemannian manifolds $\mathcal{M}_{1},\mathcal{M}_{2}$. Assume both $V_{1}$ and $V_{2}$ satisfy Assumption 1. Let $\tilde{\nu}_{1},\tilde{\nu}_{2}$ be the symmetric log-concave probability measures on $\mathbb{R}$ induced by $\nu_{1},\nu_{2}$ respectively such that $I_{\tilde{\nu}_{i}}=I_{\nu_{i}},i=1,2$. Denote $F_{i},f_{i}$ as the cdf and pdf of $\tilde{\nu}_{i}$, $i=1,2$. Define
@@
where the first infimum is taken over all smooth functions $\sigma:\mathcal{M}_{1}\to[0,1]$ such that $\int_{\mathcal{M}_{1}}\sigma(x){\text{\rm d}}\nu_{1}(x)=a$, the second infimum is taken over all smooth functions $\varphi:[0,1]\to[0,1]$ such that $\int_{0}^{1}\varphi(t){\text{\rm d}}t=a$, and the third infimum is taken over all smooth functions $g:\mathbb{R}\to\mathbb{R}$ such that $\int_{\mathbb{R}}F_{2}(g(x))f_{1}(x){\text{\rm d}}x=a$. The following proposition shows that the isoperimetric profile of $\nu_{1}\otimes\nu_{2}$ will decrease if we replace one or both of $\nu_{1},\nu_{2}$ to the corresponding $\tilde{\nu}_{1},\tilde{\nu}_{2}$. This proposition is an extension of the Gaussian version due to Bobkov [13] and the weighted Riemannian slab version due to [46] to Riemannian manifolds, and a simple proof is provided in Appendix D for completeness.
@@
holds for all smooth functions $\sigma_{i}:\mathcal{M}_{i}\to[0,1]$. Then, by repeatly applying Proposition 2, $I_{\nu_{1}\otimes\nu_{2}\otimes...\otimes\nu_{i}}\geq\Upsilon_{i},i\in\{1,2,...,n\}$.
@@
We next show that under Conjecture 2, Theorem 1 still holds for one-dimensional Euclidean space without Assumption 2. The proof of the following proposition is given in Appendix E.
@@ 1.3. Related Works
In the one-dimensional setting, Bobkov [11] established that half-lines are extremal sets in the isoperimetric inequality for log-concave probability measures. Extending to higher-dimensional structures, Bobkov and Houdré [15] further explored the isoperimetric constant and log-Sobolev inequalities in product spaces, laying a foundation for multi-dimensional isoperimetric analysis. The optimality of standard half-spaces— a core topic in isoperimetry— has been systematically studied by multiple scholars: Sudakov and Tsirel’son [61] initiated early investigations, Bobkov [11] deepened this topic by addressing the isoperimetric problem for log-concave product measures on $\mathbb{R}^{n}$ equipped with the uniform distance, and explicitly provided necessary and sufficient conditions for standard half-spaces to be extremal, and Barthe [9] investigated the extremal properties of central half-spaces (with measure $1/2$) for product probability measures on Riemannian manifolds equipped with the geodesic distance. It should be noting that the isoperimetric problem under the geodesic distance and the one under the uniform distance are significantly different. When equipped with the uniform distance, the asymptotics of the isoperimetric profile has been characterized by Bobkov in [12]. However, for the geodesic distance as in our setup, the asymptotics is still widely unknown except for special cases, e.g., the Gaussian setting. Roberto [59] offers a critical survey for a comprehensive overview of isoperimetry in product spaces. Furthermore, the isoperimery for small sets was already investigated in the literature; see, e.g., [41] for the continuous setting and [56, 53] for the discrete setting. In the discrete setting, the isoperimery for small sets is known as the small set expansion and has important applications in complexity theory [56]. It should be noted that characterizing the isoperimetric profile remains widely open, and previously so does its asymptotics.
@@
Concentration of measure has natural connections to isoperimetric inequalities. In fact, it can be seen as a variant version of isoperimetric problem concerning on sets with probability measure $1/2$ in which the boundary of a set is set to the $r$-neighbourhood of the set with $r>0$. A landmark methodological contribution in concentration of measure was made by Marton [44], who was the first to introduce information-theoretic techniques (especially transport-entropy inequalities) to the study of concentration of measure, yielding an elegant and concise proof for a dimension-free bound on concentration of measure. Talagrand extended this idea by developing a new transport-entropy inequality, adapting it to Gaussian measures and the Euclidean metric [63]; this argument has since become a standard reference, featured in prominent textbooks [38, 57, 64]. Gozlan and Léonard further refined this approach by replacing the “linear” transport-entropy inequality in Marton’s argument with a “nonlinear” version, obtaining a dimension-free sharp bound on concentration of measure—exponentially tight, as the exponent of the bound is asymptotically attained [31]. Prior to Gozlan and Léonard’s work, Alon et al. [4] are the first to characterize the sharp convergence exponent for concentration of measure, but limited to finite spaces. Gozlan [30] also used Marton’s framework to prove the equivalence between the Gaussian bound and Talagrand’s transport-entropy inequality. Additionally, Dembo [23] proposed a new class of transport-entropy inequalities, which he used them to recover and generalize several of Talagrand’s key results in [62]. Furthermore, E. Milman [45] established an equivalence between isoperimetry and concentration of measure. There are many existing works investigating isoperimetric inequalities or concentration of measure in the discrete setting, e.g., on various graphs, e.g., [43, 1, 17, 26].
Ahlswede and Zhang [3] and Yu [68] contributed to the study of isoperimetry with thick boundaries for exponentially small sets. This isoperimetric variant refines the concentration of measure phenomenon by relaxing the set size requirement to allow arbitrarily exponentially small measures, yet it behaves fundamentally differently from—and is far more intricate than—concentration of measure. This is because, in contrast to the latter, the isoperimetric variant admits no dimension-independent bounds. Ahlswede and Zhang [3] focused on finite spaces for which a key tool called the inherently typical subset lemma was developed by them. Yu [68] generalized their result to Polish spaces (complete, separable metric spaces), expanding the problem’s applicability, especially to the setting considered in this paper (i.e., the isoperimetric problem on a manifold with density), which plays a core role in our proof in the present paper.
@@
Ledoux [37, 39] proposed a semigroup approach to derive isoperimetric inequalities from hypercontractivity (which is equivalent to log-Sobolev inequalities). The semigroup approach can be used to establish isoperimetric inequalities for the whole range of volumes, but the constant obtained via this method is in general suboptimal. In contrast, the constant in our work is asymptotically optimal. Otto and Villani [54] established that log-Sobolev inequalities imply transport-entropy inequalities, bridging log-Sobolev theory with concentration of measure. Building on this, Ledoux [40] (alongside Milman [45]) used a semigroup approach to derive isoperimetric inequalities from concentration inequalities (equivalently, transport-entropy inequalities). As a nonlinear log-Sobolev inequality for the standard heat semigroups (with Lebesgue measures as stationary measures), the entropic isoperimetric inequality was investigated in [20, 24] which is in fact equivalent to the standard log-Sobolev inequality for Ornstein–Uhlenbeck semigroups (with Gaussian measures as stationary measures). Nonlinear log-Sobolev inequalities for general heat (diffusion) semigroups were investigated by F. Y. Wang [66], Bakry, Gentil, and Ledoux [7] (called the entropy-energy inequality therein), as well as Polyanskiy and Samorodnitsky [55]. Specifically, F. Y. Wang [66] showed that nonlinear log-Sobolev inequalities imply nonlinear transport-entropy inequalities, extending Otto and Villani’s linear version to the nonlinear setting. Bakry, Gentil, and Ledoux [7] and Polyanskiy and Samorodnitsky [55] applied nonlinear log-Sobolev inequalities to strengthen hypercontractive inequalities. Though nonlinear Sobolev inequalities have been studied in the literature not directly to address asymptotic isoperimetry, an exact equivalence between these two concepts does exist—one we establish in the present work. Furthermore, the equivalence between isoperimetry and Sobolev-type inequalities was investigated in [16].
1.4. Proof Idea and Paper Organization